# Real-Time Image Segmentation via Hybrid Convolutional-Transformer Architecture Search

### v1

@@ Abstract
Image segmentation is one of the most fundamental problems in computer vision and has drawn a lot of attentions due to its vast applications in image understanding and autonomous driving. However, designing effective and efficient segmentation neural architectures is a labor-intensive process that may require lots of trials by human experts. In this paper, we address the challenge of integrating multi-head self-attention into high resolution representation CNNs efficiently, by leveraging architecture search. Manually replacing convolution layers with multi-head self-attention is non-trivial due to the costly overhead in memory to maintain high resolution. By contrast, we develop a multi-target multi-branch supernet method, which not only fully utilizes the advantages of high-resolution features, but also finds the proper location for placing multi-head self-attention module. Our search algorithm is optimized towards multiple objective s (e.g., latency and mIoU) and capable of finding architectures on Pareto frontier with arbitrary number of branches in a single search. We further present a series of model via Hybrid Convolutional-Transformer Architecture Search (HyCTAS) method that searched for the best hybrid combination of light-weight convolution layers and memory-efficient self-attention layers between branches from different resolutions and fuse to high resolution for both efficiency and effectiveness. Extensive experiments demonstrate that HyCTAS outperforms previous methods on semantic segmentation task. Code and models are available at https://github.com/MarvinYu1995/HyCTAS.
I Introduction
Image segmentation predicts pixel-level annotations of different semantic categories for an image. Different from image classification, which predicts a single label for the entire image, semantic segmentation performs pixel-level labeling with a set of object categories (e.g., sky, tree, human, car) for all image pixels. Image segmentation plays a crucial role in scene understanding such as medical image analysis [1], autonomous driving [2, 3], biometrics [4, 5, 6] and image understanding [7, 6, 8]. Thanks to the enormous successes of deep neural networks, recent segmentation networks [9, 10, 11, 12, 13] have achieved impressive progresses to generate high quality results with the utilization of high resolution representation learning in network design.
@@ I Introduction
TABLE I: Comparisons with state-of-the-art search methods [14, 15, 16, 17, 18]. In which, $\rho$ and $\tau$ represent the Pearson Correlation Coefficient and the Kendall Rank Correlation Coefficient of the true accuracy (by training each of them from scratch) and estimated accuracy (predicted by the super-network), Memory Cost reports the GPU memory consumption during the search process, GPU Days presents the searching cost, Nums reports the number of the number of the discovered architectures.
@@
Recently, there are increasing interests in adapting self-attention mechanism [19] into vision tasks. Such self-attention modules can effectively model long-range dependencies and have been widely applied in natural language tasks. Although it provides an alluring alternative that solely relies on attention mechanisms, it is still far from mature to dispense convolution operators entirely because of the limitation of localization abilities due to insufficient low-level details. Hence, there is an imminent need to combine self-attention with high resolution representations.
However, naively replacing some convolution layers with self-attention layers is inefficient and non-trivial. There are two major aspects that cause the high computation cost. On one hand, the high resolution network requires to fuse features from multiple branches of different resolutions to effectively learn the semantics under various transformations. Such approaches often favor more on the prediction quality than the inference efficiency, thus resulting in heavy computation cost. For example, the latest variant of DeepLab [9] and HRNet [13] both run less than 5 FPS on a modern GPU, which is far from real-time applications. On the other hand, the network requires to maintain high-resolution representations through the whole process to generate accurate pixel-level prediction. Incorporating such resolution into self-attention layers will quadratically increase the computation cost. Due to above constraints and huge design space, manually designing effective and efficient high resolution networks with attention is a labor-intensive process that may require lots of trials from human experts.
In this paper, we propose a novel Hybrid Convolutional-Transformer Architecture Search (HyCTAS) framework to search for efficient networks with high-resolution representation and attention for superior image segmentation performance. We combine a light-weight convolution module to reduce the computation cost while preserving high-resolution information and a memory efficient self-attention module to attend long-range dependencies. We search for an efficient architecture that not only fuses features from multiple levels but also maintains high-resolution representations from a multi-branch search space with different resolutions. Finally, we deploy an efficient genetic search algorithm that is optimized towards multiple objectives (i.e. latency, mIoU) and is capable of finding series of architectures on Pareto frontier with an arbitrary number of branches from different resolutions. Experiments demonstrate that our HyCTAS outperforms previous methods on semantic segmentation task.
Our main contributions are summarized as following:
  • We design a novel searching framework incorporating with multi-branch space for high resolution representation and genetic-based multi-objective (e.g., latency and mIoU). We can obtain multiple networks covering both real-time speed and high precision within a single search at less than 5 GPU days.
  • We present a series of HyCTAS that combines a light-weight convolution module to reduce the computation cost while preserving high-resolution information and a memory efficient self-attention module to attend long-range dependencies. Such combinations enable us to search for efficient architectures with high-resolution representations with attention.
  • Extensive experiments on Cityscapes demonstrate better efficiency and effectiveness of the proposed approach over the state-of-the-art ones on semantic segmentation.
II Related Work
II-A Efficient Networks for Image Segmentation
Semantic segmentation tasks, which aim to classify each pixel into predefined classes, have greatly benefited from the advancements in deep neural networks. Initially, the Fully Convolutional Neural Network (FCN) [20] illustrated the potential of convolutional neural networks in semantic segmentation by successfully applying them to this domain. Subsequently, the DeepLab [10] architecture introduced atrous convolution, enhancing the filters’ field of view without an increase in parameters. Further improvements [9] were made by employing atrous convolution with varying atrous rates in a cascaded manner, optimizing the capture of multi-scale context. Most recently, HRNet [11, 12] was developed, which preserves high-resolution representations through the integration of high-to-low resolution convolutions and continuous fusions across parallel convolutions, leading to superior performance across various benchmarks.
While earlier segmentation methods have notably enhanced segmentation quality, the computational expense has often been overlooked, as seen in works such as those by [9, 13]. Efforts to mitigate computational costs while boosting inference speed have led to various innovations. [21] introduced a decoder network strategy to efficiently map low-resolution encoder feature maps back to the original input resolution, minimizing the necessity for high-resolution feature maintenance. [22] utilized a SqueezeNet-inspired encoder with parallel dilated convolutions to enhance efficiency. [23] achieved real-time inference speeds by reducing feature map sizes early in the network. [24] developed ICNet, integrating multi-resolution input branches rather than feature-based ones to cut computational costs. More recently, [25] proposed the Bilateral Segmentation Network (BiSeNet), which employs a novel efficient design that separates a Spatial Path to retain spatial details and a Context Path to ensure an ample receptive field.
Unlike previous works, our proposed HyCTAS does not sacrifice any high-resolution feature maps to speedup, but intelligently searches for the most efficient combination of multi-resolution branches to find the both effective and efficient architecture.
@@ II-A Efficient Networks for Image Segmentation
TABLE II: The difference between typical operation and proposed efficient operation. Flops are measured using an input of size 1$\times$256$\times$32$\times$64. The receptive field (RF) is relatively compared with the standard convolution.
@@
II-B Nerual Architecture Search
Designing effective and efficient neural architectures is a labor-intensive process that may require lots of trials by human experts. Hence, automatic neural architecture search (NAS) has emerged in recent years. Based on the searching algorithms, existing NAS methods can also be roughly divided into two categories: gradient-based differentiable methods [26, 27] and sampling-based methods [28, 29].
Recent studies have applied NAS to image segmentation tasks. [30] were pioneers in utilizing NAS for designing a network architecture comprised of stacked computation cells for semantic image segmentation.  [14] extended this by also searching for the network level structure, in addition to the cell level, to enhance segmentation. [15] introduced a method incorporating decoupled and fine-grained latency regularization to achieve a more favorable accuracy-latency trade-off. Most notably, [31] developed a framework capable of generating data-dependent routes to address scale variance in semantic representation by adapting to the scale distribution of each image.
Different from previous works, our method supports searching a combination of any number of multi-resolution branches, i.e. from 1/8 to 1/32, architectures to meet the needs of different scenarios. Compared with gradient-based differentiable methods, our proposed HyCTAS has the superiority in network capacity and higher performance with the budgeted resource consumption.
II-C Transformer in Vision Tasks
Recently, there are two major directions in adopting transformer in vision tasks. One is the pure transformer architecture, the other is the hybrid architecture which combines the CNNs and the Transformer. Vision transformer [32] was the first full self-attention based Transformer architecture without CNNs, but had limitations to transfer to downstream tasks. Meanwhile, DETR [33] employed the hybrid model architecture, which combines the CNNs with Transformer. Later, SETR [34] adapted a similar idea of vision transformer to semantic segmentation. Although transformer allows us to model full-image contextual information, it is both memory and computation intensive. Many researchers [35, 36] introduced trade-offs on speed and performance through manual designs, but have not achieved promising results yet.
@@ II-C Transformer in Vision Tasks
Figure 1: An illustration of our HyCTAS search space. It preserves the multi-resolution and multi-branch proprieties of HRNet by introducing two types of searchable components: cells and nodes, which extend HRNet with multi-module fusion ability.
@@
Figure 2: An illustration of our searchable modules. We design a memory-efficient self-attention module and a light-weight convolution module and search for the best combination.
In this paper, we also involve an efficient self-attention module as one basic component. Our proposed HyCTAS is automatically designed by setting two optimization goals (such as performance and latency) simultaneously and achieves state-of-the-art trade-off than handcrafted models.
III Proposed Approach
III-A Revisiting HRNet Architecture
@@
Learning high-resolution representations is crucial to good performance on image segmentation. HRNet achieves this by maintaining multi-branch of convolution streams in parallel to maintain high-resolution representations through the whole network, and generates reliable high-resolution representations with strong position sensitivity through repeatedly fusing the representations from multi-resolution streams.
@@
Given a supernet with $L$ layers, after an input image going through the initial stem layers, the feature resolutions are downsampled to stride $S=\{1/8,1/16,1/32\}$ and fed into multiple searchable cells $C=\{C^{1}_{s},C^{1}_{s},\dots,C^{L}_{s}\}$. Then the features are fused into searchable nodes $N=\{N^{1}_{s},N^{1}_{s},\dots,C^{L-1}_{s}\}$ consecutively, where the cell and node are alternatively stacked. The searching domain is different for cell and node. A cell could contain multiple modules such as convolution, transformer and so on, which extends the network to be capable of combining multi-module into a sub-network; Meanwhile, a node is dedicated to fuse multiple branches from different resolutions together and behaves as a binary gate.
@@
Since our architecture needs to support arbitrary branches fusion, which means the outputs of each architecture are not fixed, i.e., the number of the last feature map channels and the branch position of the final outputs are not fixed. To solve these problems, we proposed a feature aggregation head module that can adapt to arbitrary output of different architectures. The feature map of shape ($C_{*}\times H\times W$) is first adjusted in channels by a $1\times 1$ Slimmable Convolution layer [37] to the shape of ($C_{s}\times H\times W$). If the feature map size is not the max resolution in the search space, it will be bilinearly upsampled to match the shape of the other feature map ($C_{s}\times 2H\times 2W$). Then, two feature maps are concatenated and fused together with a $3\times 3$ convolution layer. We enumerate all the output combinations and set a head for each type of output $F_{\theta}$, so that there are a total of six different heads in our search space, i.e., $\theta\in\{(8),(16),(32),(8,16),(8,32),(8,16,32)\}$.
@@
Unlike previous works [14, 15] limits the network search to be one path, our HyCTAS supernet is cable of modeling sub-network with arbitrary number of branches from different resolutions, which better catches the spirit of HRNet.
III-B Combing Multiple Modules
Convolutional operations enable the extraction of visual features through weight sharing, ensuring translation invariance while maintaining sensitivity to local image regions. However, this approach does not capture relative spatial relationships, leading to a constrained receptive field.
Achieving large receptive fields is essential for tracking long-range dependencies in images, typically necessitated by employing large kernels or multiple convolutional layers, albeit at an efficiency loss. Recent advancements leverage Transformer models [19] to efficiently address the constraints of convolutional methods.
Transformers employ multi-head self-attention to model inter-pixel importance across the entire input, achieving a maximized receptive field. However, compared to convolutional methods, transformers face significant computational challenges, particularly with large feature maps, as their computational and memory demands scale quadratically with input size.
@@ III-B Combing Multiple Modules
Therefore, both operations do not meet our goal of pursuing efficiency if naively incorporated in our multi-resolution multi-branch framework for high resolution image segmentation learning. Hence, we propose to use a memory-efficient self-attention module and light-weight convolution module.
@@
Memory-efficient Self-attention Module: Instead of directly borrowing typical transformer operation like previous works [38], we only employ the self-attention layer [39] without FFN to capture the glob context while maximize the efficiency. We then use the $1\times 1$ convolutions for reducing and then increasing dimensions, leaving the self-attention module a bottleneck with smaller than input/output channels. This largely reduces the memory usages required in self-attention layer. Meanwhile, we add an extra bypassed $1\times 1$ convolution module for preserving uncompressed residue. Our memory-efficient self-attention module operation can utilize the biggest receptive field, while keeping relatively small flops and memory footprint.
@@
Light-weight Convolution Module: To maintain a large receptive field while minimizing computational costs, methods like Zoomed Convolution [15] and Dilated Convolution [9] employ bilinear sampling for input feature map resizing, inevitably leading to information loss. Addressing this, we introduce a novel, efficient convolution that incorporates a parallel $1\times 1$ convolution to preserve high-resolution details as shown in Figure 2. This approach connects input feature map directly to the zoomed convolution’s output, effectively maintaining input resolution. Our lightweight convolution module significantly reduces information loss with minimal parameter increase, ensuring both speed and accuracy.
@@
The detailed measurements of our memory-efficient self-attention module and light-weight convolution module are shown in Table 2. Meanwhile, properly combining convolutional module and self-attention module plays a significant role in the improvement of model performance and efficiency. Hence, we further deploy a search algorithm to solve this problem.
III-C Searching with Multi-objective

### v2

@@ Abstract
Image segmentation is one of the most fundamental problems in computer vision and has drawn a lot of attention due to its vast applications in image understanding and autonomous driving. However, designing effective and efficient segmentation neural architectures is a labor-intensive process that may require numerous trials by human experts. In this paper, we address the challenge of integrating multi-head self-attention into high-resolution representation CNNs efficiently by leveraging architecture search. Manually replacing convolution layers with multi-head self-attention is non-trivial due to the costly overhead in memory to maintain high resolution. By contrast, we develop a multi-target multi-branch supernet method, which not only fully utilizes the advantages of high-resolution features but also finds the proper location for placing the multi-head self-attention module. Our search algorithm is optimized towards multiple objectives (e.g., latency and mIoU) and is capable of finding architectures on the Pareto frontier with an arbitrary number of branches in a single search. We further present a series of models via the Hybrid Convolutional-Transformer Architecture Search (HyCTAS) method that searches for the best hybrid combination of light-weight convolution layers and memory-efficient self-attention layers between branches from different resolutions and fuses them to high resolution for both efficiency and effectiveness. Extensive experiments demonstrate that HyCTAS outperforms previous methods in both semantic segmentation and panoptic segmentation tasks. Code and models are available at https://github.com/MarvinYu1995/HyCTAS.
keywords:
segmentation, high-resolution, neural architecture search, attention
journal: Journal of  Templates
1 Introduction
Image segmentation predicts pixel-level annotations of different semantic categories for an image. Unlike image classification, which assigns a single label for an entire image, image segmentation requires a per-pixel label, making it a core challenge in tasks such as autonomous driving, medical image analysis, and general scene understanding. Specifically, semantic segmentation assigns each pixel to one of several predefined categories (e.g., sky, tree, human, car), instance segmentation differentiates individual object instances of the same category, and panoptic segmentation [1] unifies these two perspectives by assigning each pixel a class label and an instance identity.
Over the past decade, deep neural networks have revolutionized segmentation performance [2, 3]. In particular, high-resolution representation architectures (e.g., HRNet [4, 5]) have significantly boosted accuracy by preserving spatial details throughout the network. However, these improvements often come at the cost of high computational overhead. Manually designing a segmentation framework that retains fine spatial features, incorporates global context (e.g., via attention mechanisms), and still runs in real time is notoriously difficult.
@@ 1 Introduction
Table 1: Comparisons with state-of-the-art search methods [6, 7, 8, 9, 10]. In these comparisons, $\rho$ and $\tau$ represent the Pearson Correlation Coefficient and the Kendall Rank Correlation Coefficient of the true accuracy (by training each of them from scratch) and estimated accuracy (predicted by the super-network). Memory Cost reports the GPU memory consumption during the search process, GPU Days presents the searching cost, and Nums reports the number of the discovered architectures.
@@
Recently, transformer-based models [11, 12] have shown promise in vision tasks by modeling long-range dependencies. Yet, straightforwardly replacing convolutional layers with self-attention can be impractical for high-resolution segmentation due to quadratic complexity in both computation and memory. To achieve truly efficient, real-time segmentation, we believe it is crucial to combine:
  • 1.
    Lightweight convolutions, which preserve spatial detail while limiting computational load.
  • 2.
    Memory-efficient self-attention, which expands the receptive field and captures global dependencies without exploding memory usage.
In this paper, we propose a Hybrid Convolutional-Transformer Architecture Search (HyCTAS) framework to automatically discover powerful yet efficient segmentation networks. Rather than performing a manual, trial-and-error insertion of self-attention layers into a high-resolution CNN, HyCTAS leverages a multi-branch supernet search space and a genetic-based multi-objective optimization strategy. By jointly optimizing for latency and accuracy (mIoU), HyCTAS identifies an optimal “mix” of convolution and attention at multiple scales and resolutions. As a result, we obtain a family of architectures that lie on the Pareto frontier of speed and performance.
Our key contributions can be summarized as:
  • 1.
    Novel multi-branch search framework: We encode parallel CNN branches at multiple resolutions, each of which may contain either lightweight convolution or memory-efficient attention. This unified supernet preserves high-resolution feature streams while optimizing for real-time inference.
  • 2.
    Flexible multi-objective optimization: We employ a genetic-based algorithm to search for efficient architectures that yield a range of trade-offs between speed (latency) and segmentation accuracy (mIoU). Importantly, we can discover multiple Pareto-optimal architectures in a single search.
  • 3.
    Strong performance on multiple tasks: Extensive experiments on benchmarks such as Cityscapes, ADE20K, and COCO demonstrate that our searched models consistently outperform prior works in semantic and panoptic segmentation—while significantly reducing computational cost and increasing inference speed.
2 Related Work
2.1 Efficient Networks for Image Segmentation
Semantic segmentation has long benefited from fully convolutional networks [13], which effectively adapt classification backbones to dense prediction tasks. Subsequent refinements, such as atrous/dilated convolutions in DeepLab [14] and multi-branch high-resolution designs in HRNet [4, 5], steadily pushed performance. Meanwhile, real-time or low-complexity methods such as ICNet [15], BiSeNet [16], and Fast-SCNN [17] introduced architectural optimizations (e.g., factorized convolutions, multi-resolution branches) to reduce latency. In more recent works, lightweight backbone designs [18, 19] target embedded devices and further illustrate the trade-off between efficiency and accuracy.
Unlike most prior approaches that manually reduce resolution or remove branches to save compute, our HyCTAS approach searches across multi-resolution sub-networks, identifying the best structure automatically. This not only exploits the advantages of high-resolution representations but also finds the optimal placement of memory-efficient self-attention modules.
@@ 2.1 Efficient Networks for Image Segmentation
Table 2: The difference between a typical operation and the proposed efficient operation. FLOPs are measured using an input of size 1$\times$256$\times$32$\times$64. The receptive field (RF) is relatively compared with the standard convolution.
@@
2.2 Neural Architecture Search
Neural Architecture Search (NAS) has emerged to automate network design [20, 21]. Early NAS solutions often employed reinforcement learning or evolutionary algorithms to explore cell-based search spaces [22]. Although these methods can discover effective designs, their applicability to large-scale semantic segmentation has been limited by high computational overhead.
Recent works have extended NAS to semantic segmentation [23, 6, 7, 24], but many still rely on a single-path or single-resolution paradigm. Our HyCTAS goes further by accommodating multiple branches (with varying spatial resolutions) within one unified supernet. Moreover, our multi-objective search can discover a set of Pareto-optimal architectures—each balancing accuracy and speed differently—in a single pass.
2.3 Transformer in Vision Tasks
Transformers [25] originally showed remarkable success in NLP, and have recently gained popularity in vision tasks. Fully attentional backbones like Vision Transformer [11] and SETR [12] capture global context effectively but incur high memory costs, especially for large input resolutions. Hybrid strategies like DETR [26] combine CNNs with transformers for detection and segmentation.
@@ 2.3 Transformer in Vision Tasks
Figure 1: An illustration of our HyCTAS search space. It preserves the multi-resolution and multi-branch properties of HRNet by introducing two types of searchable components: cells and nodes, which extend HRNet with multi-module fusion ability.
@@
Figure 2: An illustration of our searchable modules. We design a memory-efficient self-attention module and a lightweight convolution module and search for the best combination.
We follow the hybrid philosophy, blending lightweight convolutions with a memory-efficient self-attention module. Crucially, we do not merely handcraft these combinations; our search algorithm determines the where and when to use attention across multiple scales, striking a better balance between accuracy and runtime efficiency.
3 Proposed Approach
3.1 Overall Workflow
Our HyCTAS framework can be summarized in four main steps:
  1. 1.
    Supernet Construction: We construct a multi-branch supernet inspired by HRNet, preserving parallel streams at various resolutions (e.g., 1/8, 1/16, 1/32). In each layer, we introduce searchable cells (for choosing lightweight convolution or memory-efficient self-attention) and searchable nodes (for fusion and skip connections).
  2. 2.
    Supernet Training: We randomly sample sub-networks from the supernet during training. To ensure coverage of multi-resolution paths, we bias the sampling probabilities so that deeper or wider sub-networks still appear sufficiently often.
  3. 3.
    Multi-Objective Search: Once the supernet weights are trained, we evaluate candidate sub-networks by measuring both segmentation accuracy (mIoU) and latency. We formulate a multi-objective problem and adopt an evolutionary algorithm (NSGA-II) to efficiently traverse the architecture space.
  4. 4.
    Architecture Selection: The search yields a set of Pareto-optimal architectures, each offering a distinct trade-off (e.g., ultra-fast inference vs. higher accuracy). We finally retrain the chosen models from scratch to confirm their performance.
3.2 Revisiting HRNet Architecture
@@
Learning high-resolution representations is crucial for good performance in image segmentation. HRNet achieves this by maintaining multi-branch convolution streams in parallel to preserve high-resolution representations throughout the whole network and generates reliable high-resolution representations with strong position sensitivity through repeatedly fusing the representations from multi-resolution streams.
@@
Given a supernet with $L$ layers, after an input image goes through the initial stem layers, the feature resolutions are downsampled to strides $S=\{1/8,1/16,1/32\}$ and fed into multiple searchable cells $C=\{C^{1}_{s},C^{2}_{s},\dots,C^{L}_{s}\}$. Then the features are fused into searchable nodes $N=\{N^{1}_{s},N^{2}_{s},\dots,N^{L-1}_{s}\}$ consecutively, where the cells and nodes are alternately stacked. The searching domain is different for cells and nodes. A cell could contain multiple modules such as convolution, transformer, and so on, which extends the network to be capable of combining multi-module into a sub-network. Meanwhile, a node is dedicated to fusing multiple branches from different resolutions together and behaves as a binary gate.
@@
Since our architecture needs to support arbitrary branch fusion, the outputs of each architecture are not fixed, i.e., the number of the last feature map channels and the branch position of the final outputs are not fixed. To solve these problems, we propose a feature aggregation head module that can adapt to arbitrary outputs of different architectures. The feature map of shape ($C_{*}\times H\times W$) is first adjusted in channels by a $1\times 1$ Slimmable Convolution layer [27] to the shape of ($C_{s}\times H\times W$). If the feature map size is not the max resolution in the search space, it will be bilinearly upsampled to match the shape of the other feature map ($C_{s}\times 2H\times 2W$). Then, the two feature maps are concatenated and fused together with a $3\times 3$ convolution layer. We enumerate all the output combinations and set a head for each type of output $F_{\theta}$, so there are a total of six different heads in our search space, i.e., $\theta\in\{(8),(16),(32),(8,16),(8,32),(8,16,32)\}$.
@@
Unlike previous works [6, 7] that limit the network search to one path, our HyCTAS supernet is capable of modeling sub-networks with arbitrary numbers of branches from different resolutions, which better captures the spirit of HRNet.
3.3 Combining Multiple Modules
Convolution facilitates the learning of visual features. With the weight-sharing mechanism, the features extracted from a convolutional layer are translation invariant. At the same time, the features are also locality-sensitive, where every operation only takes into account a local region of the image. Thus, convolutional operations fail to encode relative spatial information and result in a limited receptive field.
However, large receptive fields are often required to track long-range dependencies within an image, which in practice involves using large kernels or long sequences of convolutional layers at the cost of losing efficiency. Recent works have presented the techniques of Transformers [25] to overcome the limitations presented by convolutional operations in an efficient way.
Transformers use the concept of multi-head self-attention. Such an attention mechanism models the importance between each pixel within the whole input, which results in a maximized receptive field. Compared with convolutional operations, transformers also have the problem of large computational overhead, especially when the feature map size is very large, as its computation and memory usage are quadratic to the input size.
@@ 3.3 Combining Multiple Modules
Therefore, both operations do not meet our goal of pursuing efficiency if naively incorporated in our multi-resolution multi-branch framework for high-resolution image segmentation learning. Hence, we propose to use a memory-efficient self-attention module and a lightweight convolution module.
@@
Memory-efficient Self-attention Module: Instead of directly borrowing typical transformer operations like previous works [12, 28], we only employ the self-attention layer [25] without the FFN to capture the global context while maximizing efficiency. We then use $1\times 1$ convolutions for reducing and then increasing dimensions, leaving the self-attention module as a bottleneck with smaller than input/output channels. This largely reduces the memory usage required in the self-attention layer. Meanwhile, we add an extra bypassed $1\times 1$ convolution module for preserving uncompressed residue. Our memory-efficient self-attention module operation can utilize the largest receptive field while keeping relatively small FLOPs and memory footprint.
@@
Lightweight Convolution Module: To maintain a large receptive field while minimizing computational costs, methods like Zoomed Convolution [7] and Dilated Convolution [14] employ bilinear sampling for input feature map resizing, inevitably leading to information loss. Addressing this, we introduce a novel, efficient convolution that incorporates a parallel $1\times 1$ convolution to preserve high-resolution details, as shown in Figure 2. This approach connects the input feature map directly to the zoomed convolution’s output, effectively maintaining input resolution. Our lightweight convolution module significantly reduces information loss with minimal parameter increase, ensuring both speed and accuracy.
@@
The detailed measurements of our memory-efficient self-attention module and lightweight convolution module are shown in Table 2. Properly combining the convolutional module and the self-attention module plays a significant role in improving model performance and efficiency. Hence, we further deploy a search algorithm to solve this problem.
3.4 Searching with Multi-objective